# Fixed-Point RNNs: From Diagonal to Dense in a Few Iterations

### v1

@@ 3.1 Computing the fixed point
[Banach (1922)] Let $T(\cdot)$ be an operator with Lipschitz constant $\lambda_{T}<1$. Then, for any starting point the fixed point iteration method will converge to the fixed point of $T(\cdot)$ at rate $\mathcal{O}\left(1/\lambda_{T}\right)$.
While Banach’s theorem has been used in the context of fixed point of neural networks before (Fung et al., 2022), in this paper we are specifically focusing on recurrent models. Based on the result of Theorem 3.1, we present the following theorem for stabilizing diagonal linear RNNs:
@@
Let $f_{\theta}(x,h)$ be the diagonal linear RNN
$f_{\theta}(x,h)_{t}=\Lambda f_{\theta}(x,h)_{t-1}+\left(\mathbf{I}-\Lambda\right)\left(\mathbf{Q}\mathbf{B}_{t}x_{t}+\left(\mathbf{I}-\mathbf{Q}\right)h_{t}\right),$
where $\Lambda$ and $\mathbf{Q}$ are input-independent matrices and $\Lambda$ is diagonal. If $\Lambda$ and $\mathbf{I}-\mathbf{Q}$ are contractive (i.e. $\|\cdot\|_{2}<1$), then $f_{\theta}(x,h)$ has a Lipschitz constant $<1$ in $h$. Proof in App. A.
Theorem 3.2 provides a way to parametrize linear RNNs with input-independent transition matrices in order to ensure the latent iterated operator has Lipschitz constant $<1$: the recurrence in time is stabilized using a contractive state transition matrix $\Lambda$ and a coupled input normalization $\mathbf{I}-\Lambda$ (cf. discussion in (Zucchet & Orvieto, 2024)). Next, For stable recurrence in depth, the input transition matrix acting on $h$ is set to be contractive. This guarantees that throughout the fixed-point iteration, all sequences $h^{\ell}$ up to $h^{*}$ do not explode.
3.2 Computing the gradient
One advantage of an explicit fixed-point parameterization, such as the one derived in Theorem 3.2, lies in the gradient computation. As described by Bai et al. (2019), back-propagation through the fixed-point iteration can be avoided once a fixed-point is found. To see this, consider the Jacobian across $\ell$ iterations $\mathbf{J}_{x}^{\ell}=\frac{\partial f_{\theta}}{\partial x}(x,h^{\ell-1})$. Since $h^{\ell-1}$ depends on $x$ as well, we can recursively express $\mathbf{J}_{x}^{\ell}$ in terms of $\mathbf{J}_{x}^{\ell-1}$ and the Jacobians of a single iteration $\mathbf{J}_{x}(h)=\frac{\partial f_{\theta}}{\partial x}(x,h)$ and $\mathbf{J}_{h}=\frac{\partial f_{\theta}}{\partial h}(x,h)$ by applying the chain rule
$\mathbf{J}_{x}^{\ell}=\mathbf{J}_{x}(h^{\ell-1})+\mathbf{J}_{h^{\ell-1}}\cdot\mathbf{J}_{x}^{\ell-1}.$
The case for $\mathbf{J}_{\theta}^{*}$ works analogously. This means that the gradient w.r.t. the input $x$ and parameters $\theta$ can be computed at the fixed-point with the cost of solving $(\mathbf{I}-\mathbf{J}_{h^{*}})^{-1}$. Bai et al. (2021) and Schöne et al. (2025) approximate this inverse using the first terms of the Neumann series, which leads to a truncated backpropagation formulation or phantom gradients, incurring sequential overhead. We avoid this inversion altogether with the following workaround:
Theorem 3.3.
@@ 3.2 Computing the gradient
Let $f_{\theta}(x,h)$ be a diagonal linear RNN, with fixed-point $h^{*}$ and Lipschitz constant $<1$ in $h$. Let further $\mathcal{L}(\cdot,\cdot)$ be a loss and $y$ a target. If the Jacobians $\mathbf{J}_{x}(h)=\frac{\partial f_{\theta}}{\partial x}\left(x,h\right)$ and $\mathbf{J}_{h}=\frac{\partial f_{\theta}}{\partial h}\left(x,h\right)$ are equal, then the gradient $\nabla_{\theta}\mathcal{L}(f_{\theta}(x,h),y)$ computed at the fixed point $h=h^{*}$ will be a descent direction of $\mathcal{L}\left(f_{\theta}(x,h^{*}),y\right)$. Proof in App. B.
Similar to  (Fung et al., 2022), Theorem 3.3 shows that parametrizing $f_{\theta}(x,h)$ such that $\mathbf{J}_{x}(h)=\mathbf{J}_{h}$ guarantees optimization progress even if the gradient is computed only at the fixed-point. In practice, we observe that adhering to this condition speeds-up the convergence of the model during training.
3.3 Parameterization of the mixer $\mathbf{Q}_{t}$
In Sec. 3, we showed how to design diagonal linear recurrences that converge to a dense linear RNN via fixed-point iterations and how to train them. In this section, we turn our attention to the fixed-point dense object, and discuss a choice for $\mathbf{Q}_{t}$, where $\mathbf{A}_{t}=\mathbf{Q}^{-1}_{t}\Lambda_{t}$ as in (2), striking a balance between parameter efficiency and expressivity.
According to (Cirone et al., 2024b; Merrill et al., 2024), a key factor to increase expressivity in dense linear RNNs lies in effectively mixing information through the hidden state’s dimensions. While using a non-structured input-dependent state transition matrix would be prohibitive both computationally and in terms of required parameters with $\mathcal{O}\left(d^{3}\right)$ cost, certain structures such as circulant matrices do not improve the expressivity due to being co-diagonalizable (Cirone et al., 2024b). In the following theorem, we start with the observation that a simple low-rank parameterization could provide the necessary expressivity:
Theorem 3.4 (Informal).
While diagonal transition RNNs are confined to learning linear filters over point-wise transformations of the input path, RNNs with hidden dimension $N$ and input-dependent transition matrix of rank $\sim\log(N)$ achieve expressive universality: they can approximate any Path-to-Vector function arbitrarily well on compact domains, when $N$ is sufficiently large. Proof in App. D.
Inspired by Theorem 3.4, we start with a simple low-rank form for the mixer $\mathbf{Q}_{t}=\sum_{i=1}^{r}\alpha_{it}\mathbf{\bar{u}}_{it}\mathbf{\bar{u}}_{it}^{\top}$, where $r$ is the rank, $\alpha_{it}$ are scalar coefficients, and $\mathbf{\bar{u}}_{it}$ are unitary vectors. This structure has two benefits over a general input-dependent mixer: (1) the input-to-mixer mapping $x_{t}\rightarrow\mathbf{Q}_{t}$ requires only $\mathcal{O}(dr)$ instead of $\mathcal{O}(d^{3})$ parameters, and (2) the mixing operation has a complexity of $\mathcal{O}(dr)$ instead of $\mathcal{O}(d^{2})$.
While extremely simple and parallelizable, this parameterization requires further regularization. Following Theorem 3.2, $\mathbf{I}-\mathbf{Q}_{t}$ needs to be contractive which is satisfied if $\mathbf{Q}_{t}$ is either orthogonal or contractive as well. This can be controlled ensuring $\sum_{i}\alpha_{it}\leq 1$. Beyond that, we observe a problem of rank collapse: since the derivatives of the mixer w.r.t. the $\mathbf{u}_{it}$s are independent of each other, gradient-based optimization guides them in the same direction, resulting in a collapsed parameterization. To avoid that, one could either orthogonalize $\mathbf{u}_{it}$s or directly parametrize $\mathbf{Q}_{t}$ with orthogonal components using the following theorem:
Theorem 3.5.
(Householder factorization of unitary matrices (Urías, 2010)) A matrix $\mathbf{U}\in\mathbb{C}^{N\times N}$ is unitary if and only if for every vector $y\in\mathbb{C}^{N}$ there exists a set of Householder matrices $\{\mathbf{H}_{1},\mathbf{H}_{2},\dots,\mathbf{H}_{\ell}\}$ with $1\leq\ell\leq N$ such that $\mathbf{U}y=z\mathbf{H}_{1}\mathbf{H_{2}}\dots\mathbf{H}_{\ell}y$, where $z$ is a phase factor.
Following Theorem 3.5, we propose to parametrize $\mathbf{Q}_{t}$ as the product of a number of generalized Householder matrices
$\mathbf{Q}_{t}=\prod_{i=1}^{r}\left(\mathbf{I}-2\alpha_{it}\cdot\mathbf{\bar{u}}_{it}\mathbf{\bar{u}}_{it}^{\top}\right),$ (6)
where $10<\alpha_{it}<$ 1 and $\mathbf{\bar{u}}_{it}$s are a unit vectors. This avoids rank collapse by forcing $\mathbf{Q}_{t}$ to be full-rank, while $\mathbf{I}-\mathbf{Q}_{t}$ of rank $r$ remains contractive and stabilizes the fixed-point iteration. $\mathbf{Q}_{t}$ also has negative eigenvalues as in (Grazzi et al., 2024).
Comparision to DeltaNet.
We note the following difference with (Yang et al., 2024b): DeltaNet uses a single generalized Householder reflection as a state transition matrix within one head. While more reflections can be added by introducing zeros into the sequence (Siems et al., 2025), the interactions between channels remain constrained to a single head. Our framework allows for a principled parametrization of dense linear RNNs as fixed-points of diagonal linear RNNs with various matrix structures $\mathbf{Q}_{t}$ to mix across all available channels.
4 Memory in Fixed-Point RNNs
In this section, we identify conditions which improve the memory capabilities of Fixed-Point RNNs. We start by improving the propagation of information through time and then evaluate Fixed-Point RNNs on the copy task introduced by Jelassi et al. (2024) as a fundamental benchmark for memory.
4.1 Accelerated state propagation
An important factor for better memory in a recurrent model is the free flow of information through time. However, in the fixed-point mechanism introduced in (4), information propagates at the same speed in-depth and time. As a result, it may take $T$ iterations for the model to propagate the hidden state across a sequence of length $T$. Consequently, the number of fixed-point iterations is directly coupled to the sequence length, resulting in a quadratic runtime $O(T^{2})$ akin to softmax-attention. Luckily, if the fixed-point $h^{*}$ is unique, any method would converge to it. Therefore, the fixed-point iteration can be accelerated by iterating through time once per fixed-point step, i.e. by replacing $h_{t-1}^{{\ell}-1}$ with $h_{t-1}^{{\ell}}$ as in
$h_{t}^{\ell}=\lambda_{t}\odot h_{t-1}^{{\ell}}+\mathbf{Q}_{t}\mathbf{B}_{t}x_{t}+(\mathbf{I}-\mathbf{Q}_{t})h_{t}^{\ell-1}.$ (7)
Here, $\Lambda_{t}$ is replaced by its diagonal $\lambda_{t}$, and $\odot$ is the element-wise product. This allows computing one fixed-point step in $\mathcal{O}(T)$ with existing implementations based on the parallel scan (Martin & Cundy, 2018), while also improving the information propagation through time when computing the fixed-point. From now on, we rely on this form of computing the fixed-point instead of the formulation in (4).
4.2 Introducing matrix hidden states
@@
Memory capacity is another important consideration in RNNs. In preliminary experiments, we notice a clear gap between the performance of a Fixed-Point RNN and Mamba (Dao & Gu, 2024) in terms of copying ability. We attribute this difference in performance to Mamba’s state-expansion which endows it with matrix hidden states similar to linear attention, DeltaNet, or mLSTM  (Katharopoulos et al., 2020; Schlag et al., 2021b; Beck et al., 2024). In simple terms, these models use an outer product of an input-dependent vector $\mathbf{B}_{t}\in\mathbb{R}^{d_{\text{state}}}$ and the input $x_{t}\in\mathbb{R}^{d_{\text{inner}}}$ as a matrix-valued input to a recurrence $h_{t}=\lambda_{t}\odot h_{t-1}+\mathbf{B}_{t}x_{t}^{\top}$ with matrix-valued hidden state and transition gate $h_{t},\lambda_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$. The hidden state is then contracted with another input-dependent vector $\mathbf{C}_{t}\in\mathbb{R}^{d_{\text{state}}}$ to get the output $y_{t}^{\top}=\mathbf{C}_{t}^{\top}h_{t}\in\mathbb{R}^{d_{\text{inner}}}$.
@@
The matrix-valued recurrence introduces some challenges to our fixed-point framework. Specifically, in order to mix all the channels over the entirety of the state elements, the mixer $\mathbf{Q}_{t}$ has to be a fourth-order tensor $\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}\times d_{\text{state}}\times d_{\text{inner}}}$ in
$h_{t}^{\ell}=\lambda_{t}\odot h_{t-1}^{\ell}+\mathbf{Q}_{t}\otimes\mathbf{B}_{t}x_{t}^{\top}+\left(\mathbf{I}_{d^{4}}-\mathbf{Q}_{t}\right)\otimes h_{t}^{\ell-1},$ (8)
@@
where, $\otimes$ denotes the fourth-order tensor product and $\mathbf{I}_{d^{4}}$ a fourth-order identity tensor of the same shape as $\mathbf{Q}_{t}$. Certainly computing the fixed-point introduced in (8) will be very challenging both in terms of computation and memory. As we will confirm in Section 5, one solution is to pass the contracted output $y_{t}$ between fixed-point iterations
$h_{t}^{\ell}=\lambda_{t}\odot h_{t-1}^{\ell}+\mathbf{B}_{t}\left(\mathbf{Q}_{t}x_{t}\right)^{\top}+\mathbf{B}_{t}\left(\left(\mathbf{I}-\mathbf{Q}_{t}\right)y_{t}^{\ell-1}\right)^{\top}.$ (9)
@@
This implicitly factorizes the mixer $\mathbf{Q}_{t}$ into separately mixing along dimension $d_{\text{inner}}$ which is used for better expressivity, and dimension $d_{\text{state}}$ which is used for better memory. For consistency, we continue to denote the hidden state as $h_{t}^{\ell}$ until we define the Bedouin model in Sec. 5.
4.3 Dependence on $h_{t-1}$ in practice
Unfortunately, even the Fixed-Point RNN with input-dependent parameters and matrix state akin to Mamba-1 is outperformed by Mamba-2 or DeltaNet (Dao & Gu, 2024; Yang et al., 2024b) on the copy task. Inspired by the short convolution in Mamba, we investigate the effect of augmenting the input-dependence of parameters $\lambda_{t}^{\ell}$, $\mathbf{B}_{t}^{\ell}$, $\mathbf{C}_{t}^{\ell}$, and $\mathbf{Q}_{t}^{\ell}$ at iteration $\ell$ with a shifted hidden-state dependence. In practice, this means that these are linear functions of $x_{t}$ as well as the shifted previous iterate (layer) $h_{t-1}^{\ell-1}$. We refer the reader to Sec. 5.3 for the exact formulation of the dependency.
Dependence on $h_{t-1}^{\ell-1}$ Test Accuracy
$\lambda_{t}$ $\mathbf{Q}_{t}$ $\mathbf{B}_{t}$ $\mathbf{C}_{t}$

### v2

@@ 3.2 The Fixed-Point Iteration
Let $f_{\theta}(x,h)$ be the diagonal linear RNN with input-independent $\mathbf{\Lambda}$ and $\mathbf{Q}$ $f_{\theta}:(x,h)\mapsto h^{\prime},$ $h^{\prime}_{t}=\mathbf{\Lambda}h^{\prime}_{t-1}+\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{Q}\mathbf{B}_{t}x_{t}+\left(\mathbf{I}-\mathbf{Q}\right)h_{t}\right).$ (7) If $||\mathbf{\Lambda}||_{2}<1$ and $||\mathbf{I}-\mathbf{Q}||_{2}<1$, then $f_{\theta}(x,h)$ has a Lipschitz constant $<1$ in $h$. Proof in App. B.1.
Intuitively, Thm. 3.1 states two conditions for stable parametrization of an implicitly dense RNN $F_{\theta}$: (1) the recurrence in time needs to be coupled with input normalization and contractive (i.e. $\|\mathbf{\Lambda}\|_{2}<1$). (2) The recurrence in depth acting on $h$, i.e. $(\mathbf{I}-\mathbf{Q}_{t})$, needs to be contractive. Together, this guarantees that all sequences $h^{\ell}$ up to $h^{*}$ throughout the fixed-point iteration do not explode without any explicit assumptions on the spectral radius on $\mathbf{A}$ (Arjovsky et al., 2016).
3.3 Parametrization of $\mathbf{Q}_{t}$ and $\mathbf{\Lambda}_{t}$
To satisfy the assumptions required for expressivity in (Cirone et al., 2024b), the implicit transition matrix $\mathbf{A}_{t}$ and therefore $\mathbf{\Lambda}_{t}$ and $\mathbf{Q}_{t}$ need to be input-controlled (i.e. selective), which could be realized through a linear mapping of the input, i.e. $\mathbf{Q}_{t}=\mathcal{M}(x_{t}):=\text{reshape}(\mathbf{W_{Q}}x_{t})$. However, this presents two challenges: how can stability be guaranteed (c.f. Thm. 3.1) and excessive computational cost due to the $O(d^{3})$ parameters of $\mathbf{W_{Q}}$ be avoided? A straight-forward solution lies in structured matrix representations for both the diagonal transition matrix $\mathbf{\Lambda}_{t}$ and the channel mixer $\mathbf{Q}_{t}$.
Inspired by Helfrich et al. (2018), we aim for $\mathbf{Q}_{t}$ to be approximately norm-preserving and $\mathbf{\Lambda}_{t}$ to control the eigenvalue scale using a parametrization akin to Mamba or Griffin (Gu & Dao, 2023; De et al., 2024) and normalization $(\mathbf{I}-\mathbf{\Lambda}_{t})$. For the channel mixers $\mathbf{Q}_{t}$, we consider the structures:
  • Diagonal Plus Low Rank (DPLR): $\mathbf{Q}_{t}=\mathcal{M}(x_{t}):=\left(\mathbf{I}-\sum_{i=1}^{r}\alpha_{it}\cdot\mathbf{\bar{u}}_{it}\mathbf{\bar{u}}_{it}^{\top}\right)$, for rank $r$.
  • Householder Reflections (H): $\mathbf{Q}_{t}=\mathcal{M}(x_{t}):=\prod_{i=1}^{r}\left(\mathbf{I}-\alpha_{it}\cdot\mathbf{\bar{u}}_{it}\mathbf{\bar{u}}_{it}^{\top}\right)$, for $r$ reflections.
  • Kronecker (K): $\mathbf{Q}_{t}=\mathcal{M}(x_{t}):=\mathbf{I}-(\mathbf{\bar{K}}_{t}^{1}\otimes\mathbf{\bar{K}}_{t}^{2})$, where $\otimes$ denotes the Kronecker product.
This allows to reduce the size of the input-dependent parameters $\alpha_{it}$, $\mathbf{\bar{u}}_{it}$, and $\mathbf{\bar{K}}_{t}^{i}$ to $O(d)$, and consequently reduce the size of the linear map $\mathbf{W_{Q}}$ to $O(d^{2}r)$ and $O(d^{2})$. In order to guarantee stability, the condition $||\mathbf{I}-\mathbf{Q}||_{2}<1$ can be enforced by scaling $\alpha_{it}$, $\mathbf{\bar{u}}_{it}$, and $\mathbf{\bar{K}}_{t}^{i}$ appropriately. For more details about the channel mixer variants and how to scale them to negative eigenvalues, please refer to App. C. In Fig. 3(b), we compare different channel mixer variants and observe that the Kronecker structure seems to be most appropriate the state-tracking task $A_{5}$.
3.4 Algorithmic Implications
Recall from Sec. 2 that an explicitly parametrized dense selective RNN can only be parallelized under strict assumptions on its structure and runs otherwise in $O(T)$ sequential steps. However, a parallelizable structure is given by the element-wise, diagonal transition $\mathbf{\Lambda}_{t}$ of a diagonal RNN (Martin & Cundy, 2018). Since such a diagonal RNN is called $\ell^{*}$-times as a subroutine of the fixed-point iteration in Eq. 6, a fixed-point RNN runs in $O(\ell^{*}\cdot\log{T})$ sequential steps. This means that the implicit parametrization –as opposed to explicit or non-linear parametrizations– allows to decouple the number of sequential steps $\ell^{*}$ from the sequence length $T$ itself, and trade parallelism for expressivity.
This insight suggests an opportunity to introduce a non-linear computation for every sequential step, like in classical RNNs. Concretely, we investigate channel mixers $\mathcal{M}(x_{t}+\smash{h_{t-1}^{\ell-1}})$ which are a function of both the input $x_{t}$ and the hidden state $\smash{h_{t-1}^{\ell-1}}$ from the previous iteration (in both time and depth) without degrading parallelizability. In Fig. 3(b), we compare channel mixers with and without hidden state dependence and observe that this indeed improves sequence length generalization.
Summarizing the results so far, we arrive at an updated recurrence with hidden state dependence:
$h_{t}^{\ell}=\lambda_{t}^{\ell}\odot h_{t-1}^{{\ell}}+(\mathbf{1}-\lambda_{t}^{\ell})\odot\big{(}\mathbf{Q}_{t}^{\ell}\mathbf{B}_{t}^{\ell}x_{t}+(\mathbf{I}-\mathbf{Q}_{t}^{\ell})h_{t}^{\ell-1}\big{)},$ (8)
where we use $\odot$ to highlight the parallelizability of the element-wise product. We would like to note that due to the normalization $(\mathbf{I}-\mathbf{\Lambda}_{t})$, the corresponding dense RNN $F_{\theta}$ is not explicitly representable anymore. Furthermore, for the time-varying parametrization in Eq. 8, the convergence guarantees may be weaker and solutions $h^{*}$ could be non-unique due to the hidden state dependence. In practice, we iterate until $\frac{||h^{\ell}-h^{\ell-1}||_{\infty}}{||h^{\ell}||_{\infty}}<0.1$ and observe that the conditions of Thm. 3.1 are strong enough to reach convergence as evidenced by Fig. 3(c). Interestingly, the model navigates the parallelism tradeoff (Merrill & Sabharwal, 2023b) and adaptively increases its sequential computation for harder tasks.
3.5 Optimizing Fixed-Point RNNs
One advantage of converging to a fixed-point as opposed to general layer looping lies in model training. Since the gradient with respect to $h^{0}$ is not needed, implicit differentiation can be used to avoid storing and backpropagating through the computational graph of the fixed-point iteration, as discussed by Liao et al. (2018), Bai et al. (2019), and in App. B.2. In practice, truncated backpropagation of the last $k$ iterations suffices to approximate the gradient through the full iteration $\mathbf{J}_{x}^{*}\approx\mathbf{J}_{x}(h^{\ell^{*}-k})\cdot\ldots\cdot\mathbf{J}_{x}(h^{\ell^{*}})$. For Fixed-Point RNNs we observe that computing the gradient only at the fixed-point ($k=0$), is enough to stabilize training. This means that compared to a single diagonal RNN layer, Fixed-Point RNNs incur no memory overhead and only sequential overhead in the forward pass but not in the backward pass.
We hypothesize that this is possible because $f_{\theta}(x,h)$ is a mostly linear object as opposed to multi-layer implicit models such as (Schöne et al., 2025). Furthermore, we observe that hidden state dependence $\mathcal{M}(x_{t}+h_{t-1}^{\ell-1})$ particularly helps with gradient-based optimization. We credit this to the symmetry between the gradients w.r.t. $x$ and $h$, and formalize this in the following theorem:
@@ 3.5 Optimizing Fixed-Point RNNs
Let $f_{\theta}(x,h)$ have Lipschitz constant $<1$ and fixed-point $h^{*}$. If the Jacobians $\frac{\partial f_{\theta}}{\partial x}(x,h)$ and $\frac{\partial f_{\theta}}{\partial h}(x,h)$ are equal, then the gradient $\nabla_{\theta}\mathcal{L}\big{(}f_{\theta}(x,h),y\big{)}$ of the loss $\mathcal{L}(\cdot,y)$ for a target $y$ at the fixed point $h=h^{*}$ will be a descent direction of $\mathcal{L}\big{(}F_{\theta}(x),y\big{)}$. Proof in App. B.3.
4 Fixed-Point Mamba
Figure 4: Length generalization on $A_{5}$ (a, c) and $S_{5}$ (b, d) beyond the train sequence length $16$ (pink line). We compare a 1-layer FP-Mamba with mixer variants $\mathbf{Q}_{t}$ to baselines with 2 layers.
In the previous section we introduced the FP-RNN framework on a small RNN with vector hidden state. Now, we extend it to modern matrix state RNNs in Sec. 4.1 and parametrize a dense variant of Mamba (Gu & Dao, 2023) in Sec. 4.2. A detailed description of the architecture is available in App. C.2. We compare the architecture to the baselines Mamba (Gu & Dao, 2023), Mamba-2 (Dao & Gu, 2024), Gated DeltaNet (Yang et al., 2025), and LSTM (Hochreiter & Schmidhuber, 1997) on the copy task introduced by Jelassi et al. (2024) in Sec. 4.3 and state-tracking introduced by Merrill & Sabharwal (2023b) in Sec. 4.4. In order to keep the number of layers at the same order of magnitude, we use two layers (2L) for the diagonal linear RNN baselines and one layer (1L) for FP-Mamba and LSTM. Finally, we discuss the required number of fixed-point iterations in the context of state-tracking and language modeling in Sec. 4.5.
4.1 Introducing Matrix States
@@
Memory capacity is an important consideration in RNNs. In preliminary experiments, we notice a clear gap between the performance of a Fixed-Point RNNs and Mamba in terms of copying ability. We attribute this difference in performance to Mamba’s state-expansion which endows it with matrix hidden states similar to linear attention, DeltaNet, or mLSTM  (Katharopoulos et al., 2020; Schlag et al., 2021; Beck et al., 2024). In simple terms, these models use an outer product of an input-dependent vector $\mathbf{B}_{t}\in\mathbb{R}^{d_{\text{state}}}$ (i.e. the key) and the input vector $x_{t}\in\mathbb{R}^{d_{\text{inner}}}$ (i.e. the value) as an input to a matrix-valued recurrence with hidden state and transition gate $\mathbf{h}_{t},\boldsymbol{\lambda}_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$. The hidden state is then contracted with another input-dependent vector $\mathbf{C}_{t}\in\mathbb{R}^{d_{\text{state}}}$ (i.e. the query) to get the output $y_{t}^{\top}=\mathbf{C}_{t}^{\top}\mathbf{h}_{t}\in\mathbb{R}^{d_{\text{inner}}}$:
$\mathbf{h}_{t}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}+\mathbf{B}_{t}x_{t}^{\top},$ (9)
@@
This matrix-valued recurrence introduces some challenges to our fixed-point framework. Specifically, in order to mix all the channels over the entirety of the state elements, the mixer $\mathbf{Q}_{t}$ has to be a fourth-order tensor $\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}\times d_{\text{state}}\times d_{\text{inner}}}$ in
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{Q}_{t}\otimes\mathbf{B}_{t}x_{t}^{\top}+\left(\mathbf{I}_{d^{4}}-\mathbf{Q}_{t}\right)\otimes\mathbf{h}_{t}^{\ell-1},$ (10)
@@
where, $\otimes$ denotes the fourth-order tensor product and $\mathbf{I}_{d^{4}}$ a fourth-order identity tensor of the same shape as $\mathbf{Q}_{t}$. Certainly computing the fixed-point introduced in Eq. 10 is very challenging both in terms of computation and memory. As we will confirm in Sec. 4.2, one solution is to pass the contracted output $y_{t}$ between fixed-point iterations
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{B}_{t}\left(\mathbf{Q}_{t}x_{t}\right)^{\top}+\mathbf{B}_{t}\left(\left(\mathbf{I}-\mathbf{Q}_{t}\right)y_{t}^{\ell-1}\right)^{\top}.$ (11)
@@
This implicitly factorizes the mixer $\mathbf{Q}_{t}$ into separately mixing along dimension $d_{\text{inner}}$ which is used for better expressivity, and dimension $d_{\text{state}}$ which is used for better memory capacity.
4.2 FP-Mamba Iteration
Let us apply the the fixed-point RNN framework to the Mamba parametrization. We represent the hidden state as $\mathbf{h}_{t}^{\ell}$, where $t$ is the token index (i.e., indexing over the sequence dimension), and $\ell$ is the fixed-point iteration index (i.e., indexing over the depth dimension). The same notation is used for other variables to emphasize when they depend on the input and hidden state of the current iteration. We propose the following iteration to adapt Mamba to the fixed-point mechanism for matrix state RNNs in Eq. 11:
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\mathbf{Q}_{t}^{\ell}x_{t}\right)^{\top}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\left(\mathbf{I}-\mathbf{Q}_{t}^{\ell}\right)y_{t}^{\ell-1}\right)^{\top},$ ${y_{t}^{\ell}}^{\top}=({\mathbf{\bar{C}}_{t}^{\ell}})^{\top}\mathbf{h}_{t}^{\ell}.$ (12)
L2-normalizing $\mathbf{\bar{B}}_{t}^{\ell}$ and $\mathbf{\bar{C}}_{t}^{\ell}$ allows to limit the Lipschitz constant according to Theorem 3.1. Furthermore, we replace the normalization term $(\mathbf{1}-\boldsymbol{\lambda}_{t})$ with Mamba’s normalization term $\Delta_{t}$. Expanding $y_{t}^{\ell-1}$ yields the recurrence on the matrix state
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{\bar{B}}_{t}^{\ell}(\Delta_{t}\mathbf{Q}_{t}^{\ell}x_{t})^{\top}+\mathbf{\bar{B}}_{t}^{\ell}(\mathbf{\bar{C}}_{t}^{\ell-1})^{\top}\mathbf{h}_{t}^{\ell-1}(\mathbf{I}-\mathbf{Q}_{t}^{\ell})^{\top}\Delta_{t},$
where the last term nicely illustrates the two components which mix the channels of the hidden states: the low-rank matrix $\mathbf{\bar{B}}_{t}^{\ell}(\mathbf{\bar{C}}_{t}^{\ell-1})^{\top}$ mixes over the dimension $d_{\text{state}}$, while $(\mathbf{I}-\mathbf{Q}_{t}^{\ell})^{\top}$ mixes over the dimension $d_{\text{inner}}$. This factorization significantly simplifies the fourth-order tensor mixer formulation introduced in Eq. 10 and performs well in practice.
Finally, Eq. 12 can be computed as Mamba with an adjusted input $\tilde{x}_{t}^{\ell}=\mathbf{Q}_{t}^{\ell}\left(x_{t}-y_{t}^{\ell-1}\right)+y_{t}^{\ell-1}$,
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\tilde{x}_{t}^{\ell}\right)^{\top}.$ (13)
Dependence on $y_{t-1}^{\ell-1}$ Test Accuracy
$\boldsymbol{\lambda}_{t}$ $\mathbf{Q}_{t}$ $\mathbf{B}_{t}$ $\mathbf{C}_{t}$